# The Central Role of the Loss Function in Reinforcement Learning

### v1

@@ 4.7 Fitted Q-Iteration Algorithms for hybrid RL
Importantly, we see that simply changing the loss from $\ell_{\operatorname{sq}}$ to $\ell_{\operatorname{bce}}$ again leads to improved first-order bounds, which again supports our earlier observations. Compared with our prior results, the main advantage of Thm. 12 is computational: it bounds the sub-optimality of a computationally efficient algorithm FQI, which much more closely resembles deep RL algorithms such as DQN [36]. From a statistical perspective, the hybrid RL bound is actually worse than either pure online or offline bounds, since it takes the form:
@@ 5 Discussion and Conclusion
From the one-step CSC to online, offline and hybrid RL, we see time and time again that the loss function plays a central role in the adaptivity and efficiency of decision making algorithms. The classical squared loss always converges at a slow $\widetilde{\mathcal{O}}(1/\sqrt{n})$ rate and cannot adapt to easier problem instances with heteroskedasticity. The bce loss can serve as a drop-in improvement that yields first-order bounds with a much faster $\widetilde{\mathcal{O}}(1/n)$ rate when the optimal cost is small. Switching from conditional-mean learning to conditional-distribution learning, the mle loss can tighten the bounds further with a second-order guarantee, that is bounds that converge at a $\widetilde{\mathcal{O}}(1/n)$ rate in near-deterministic settings even if the optimal cost is large. Crucially, these gaps in performance are not merely theoretical as they have been observed many times by the deep RL community [15, 7, 21, 4, 33]. The tools and principles outlined in this paper are very general and can be applied to a wide range of problems including imitation learning [17], model-based RL [48], and risk-sensitive RL [49]. Thus, we hope to have not only demonstrated clearly that loss function choice is important in RL, but also to inspire the reader to seek out opportunities for better loss functions to improve any decision-making algorithm.
@@ 6 Appendix
Claim 2 (Pigeonhole): For any $\varepsilon_{0}$ and any sequence $p^{(1)},\dots,p^{(N)}\in\mathcal{P}$, there exists $j\leq N$ s.t. $N(p^{(j)},p^{(<j)},\varepsilon_{0})\geq\lfloor\frac{(N-1)}{\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})}\rfloor$. Recall that if $p^{(1)},\dots,p^{(L)}\subset\mathcal{P}$ satisfies for all $j\in[L]$, $p^{(j)}$ is $\varepsilon_{0}$-independent of $p^{(<j)}$, then $L\leq\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})$ by definition. To prove the claim, we maintain $J:=\lfloor\frac{(k-1)}{\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})}\rfloor$ disjoint sequences $S^{(1)},\dots,S^{(J)}\subset p^{(<k)}$ s.t. each $S^{(i)}$ has the property that each element is $\varepsilon$-independent of its precedessors. We initialize $S^{(1)}=\cdots=S^{(J)}=\emptyset$ and iteratively add elements $p^{(1)},\dots,p^{(N)}$ until $p^{(j)}$ is $\varepsilon_{0}$-dependent on all these disjoint subsequences, at which point the claim is proven. If there exists a subsequence which $p^{(j)}$ is $\varepsilon_{0}$-independent of, we add $p^{(j)}$ to that subsequence, which preserves the invariant condition. This process indeed terminates since otherwise one subsequence would have more elements than $\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})$, a contradiction.

### v2

@@ 4.7 Fitted Q-Iteration Algorithms for hybrid RL
Importantly, we see that simply changing the loss from $\ell_{\operatorname{sq}}$ to $\ell_{\operatorname{bce}}$ again leads to improved first-order bounds, which again supports our earlier observations. Compared with our prior results, the main advantage of Thm. 12 is computational: it bounds the sub-optimality of a computationally efficient algorithm FQI, which much more closely resembles deep RL algorithms such as DQN [48]. From a statistical perspective, the hybrid RL bound is actually worse than either pure online or offline bounds, since it takes the form of:
@@ 5 Discussion and Conclusion
From the one-step CSC to online, offline and hybrid RL, we see time and time again that the loss function plays a central role in the adaptivity and efficiency of decision making algorithms. The classical squared loss always converges at a slow $\widetilde{\mathcal{O}}(1/\sqrt{n})$ rate and cannot adapt to easier problem instances with heteroskedasticity. The bce loss can serve as a drop-in improvement that yields first-order bounds with a much faster $\widetilde{\mathcal{O}}(1/n)$ rate when the optimal cost is small. Switching from conditional-mean learning to conditional-distribution learning, the mle loss can tighten the bounds further with a second-order guarantee, that is bounds that converge at a $\widetilde{\mathcal{O}}(1/n)$ rate in near-deterministic settings even if the optimal cost is large. Crucially, these gaps in performance are not merely theoretical as they have been observed many times by the deep RL community [19, 10, 29, 7, 45]. The theory outlined herein is very general and can be applied to a wide range of settings including imitation learning [21], model-based [24, 63], risk-sensitive RL [60, 64], and robust bandits [36], and RL [12]. Moreover, the principles herein can lead to better algorithms for real-world applications such as pre-training and post-training large language models [27, 1, 65], learning query optimizers [40, 66] and many more. We hope to have not only clearly demonstrated that the loss function choice is important in RL, but also to inspire the reader to seek out opportunities for better loss functions to improve their decision-making algorithms.
@@ 6 Appendix
Claim 2 (Pigeonhole): For any $\varepsilon_{0}$ and any sequence $p^{(1)},\dots,p^{(N)}\in\mathcal{P}$, there exists $j\leq N$ such that $N(p^{(j)},p^{(<j)},\varepsilon_{0})\geq\lfloor\frac{(N-1)}{\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})}\rfloor$. Recall that if $p^{(1)},\dots,p^{(L)}\subset\mathcal{P}$ satisfies for all $j\in[L]$, $p^{(j)}$ is $\varepsilon_{0}$-independent of $p^{(<j)}$, then $L\leq\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})$ by definition. To prove the claim, we maintain $J:=\lfloor\frac{(k-1)}{\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})}\rfloor$ disjoint sequences $S^{(1)},\dots,S^{(J)}\subset p^{(<k)}$ s.t. each $S^{(i)}$ has the property that each element is $\varepsilon$-independent of its precedessors. We initialize $S^{(1)}=\cdots=S^{(J)}=\emptyset$ and iteratively add elements $p^{(1)},\dots,p^{(N)}$ until $p^{(j)}$ is $\varepsilon_{0}$-dependent on all these disjoint subsequences, at which point the claim is proven. If there exists a subsequence which $p^{(j)}$ is $\varepsilon_{0}$-independent of, we add $p^{(j)}$ to that subsequence, which preserves the invariant condition. This process indeed terminates since otherwise one subsequence would have more elements than $\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})$, a contradiction.